# Kishu: Time-Traveling for Computational Notebooks

### v3

@@ 2.2. Characteristics of Notebook Workloads
Notebook workflows often follow a loop of users sequentially [^2] Existing notebook systems, e.g., Jupyter, do not support concurrent cell executions. writing and running cell code, then observing output (Rule et al., 2018). The notebook kernel can be inactive mid-loop—$\sim$10 seconds of think time (Eichmann et al., 2020) when users decide what cells to write and run next. Notebook systems use think time for computations (e.g., speculative materialization (Xin et al., 2021)). Kishu can also leverage think time for incremental checkpointing.
@@
Python functions in notebooks can exhibit complex access patterns (e.g., accessing non-parameter variables), making it difficult and/or costly to analyze changes. Hence, some works use a rule-based approach to hard-code effects of common functions (e.g., print not modifying the namespace) (Macke et al., 2020); however, as data scientists also commonly implement and import custom functions into notebooks (Rule et al., 2018), an efficient approach that handles arbitrary functions is highly desired.
@@
Users often run cells incrementally, each consisting of only a few lines of code accessing a small number of variables, for fast iteration (Kery et al., 2018). Accordingly, we observe in on of our test notebooks, Sklearn (LABEL:fig:background_workload_characteristics, top), that 40/44 cells access ¡10% of the state data. For these workloads, Kishu’s incremental checkpointing can offer larger benefits.
@@
Notebook providers often limit session memory consumption (Google, 2024; Kaggle, 2024). Hence, in addition to creating new data, users also modify/delete existing data to conserve memory (e.g., after they have been used to generate relevant cell outputs). This is seen in Sklearn (LABEL:fig:background_workload_characteristics, bottom), where updated data is, on average, split 45%:55% between creations and modifications. Kishu must efficiently undo these in-place updates during checkout.
@@ 2.3. Enabling Time Traveling
Tools such as CRIU (CRIU, 2023) can snapshot notebook process memory to save session state data. (Incremental Checkpointing) Subsequent snapshots can be made incrementally w.r.t. previous snapshots storing only dirty memory pages. However, memory-page granularity is too coarse for incrementally checkpointing notebook workloads as Python data structures (e.g., lists) can be fragmented, leading to simple operations (e.g., in-place sorting a list) causing multiple dirty pages and high checkpoint costs (§ 7.3). (Complete Checkout) Memory snapshots must be entirely loaded to restore a state and additionally require killing the existing notebook process (to avoid PID conflicts) before restoration, which is not seamless and incurs high data loading costs (§ 7.5). Furthermore, OS-level snapshotting is limited to single processes, hence fail on notebooks utilizing multiple/remote processors (§ 7.3).
Application-level Incremental Checkpoint and Checkout (Ours)
We can utilize (1) application-level information to detect state deltas at a finer granularity compared to memory snapshots and (2) existing data within the kernel for incremental checkout to address the drawbacks of OS-level snapshotting: (Incremental and Generalized Checkpointing) We track state deltas at a more logical Co-variable granularity (described and motivated in the next session) by tracking references in the notebook for incremental checkpointing. For generalizability, we utilize an object’s reduction (i.e., __reduce__ (Foundation, 2023f)) as storage instructions to checkpoint multiprocessing and off-CPU workloads (§ 6.1). (Incremental Checkout) As we have direct access to the notebook kernel at the application level, we can incrementally checkout by only updating data in the kernel that differ between the current and target states after computing said difference (e.g.., via state snapshots (Bernstein and Goodman, 1983)) (§ 5.2).
@@ 2.4. Tracking State Delta for Time Traveling
Some existing notebook systems (Macke et al., 2020; Koop and Patel, 2017; Li et al., 2023a) track session state evolution via provenance-based code analysis (e.g., via ASTs (Foundation, 2023a)) at a variable-level granularity. Pure static code analysis requirew conservativeness on identifying changed data w.r.t. control flows (e.g., if (x<1):) and external function calls (§ 2.2) causing false positives (e.g., by assuming an untaken branch as taken) and large deltas; hence, these systems augment code analysis with varying degrees of live instrumentation at cell runtime (e.g., resolving the value of x when evaluating the aforementioned if statement) (Macke et al., 2020), which can potentially result in high overhead (e.g., repeated resolutions in looping control flows, § 7.6).
@@
In order to avoid potential efficiency issues caused by the runtime resolution of provenance tracking, we propose to perform live object comparison (i.e., comparing data before and after execution) only after each cell execution to track per-cell updates. Our intuition is that while doing so at variable-level granularity (i.e., like existing provenance trackers) by comparing all objects in the state can be expensive (§ 7.6), it is also unnecessary as storing/loading individual variables (e.g., with variable-level KV-stores (shove, 2024; 0xnurl, 2024)) during checkpointing/checkout risks breaking shared references (Li et al., 2023a). We hence track updates at a coarser Co-variable level—connected components of objects (w.r.t. pointer references): we can reason based on access patterns which Co-variables were possibly/definitely not updated by each cell to limit object comparisons, and correctly store/load Co-variables as if they are independent data tables. LABEL:fig:background_variable_blob depicts our idea: {ser,obj} is a Co-variable (red) as the reachable objects of the variables overlap (&ser[1] = &obj.foo). {df} is another Co-variable (blue), and users cannot reach objects under df from objects under ser or obj via references. Notably, Co-variables are the minimum granularity which data in the state can be stored/loaded without risking breaking shared references. § 4 formally describes Co-variables and how we correctly and efficiently capture state delta at this granularity.
@@
Suppose a data analyst is performing text mining. They load the corpus (Cell 1), define category lists (Cell 2), and sort texts by sentiment into the lists (Cell 3). The analyst incrementally checkpoints the state after each cell execution. (1) Incremental Checkpointing: The analyst tests a mapping to clean the contained text in the lists on sad_ls (Cell 4, blue). Due to its interleaved construction (with other lists), the list sad_ls is fragmented; incrementally checkpointing the state at memory page granularity for Cell 4 (w.r.t. Cell 3) results in all pages overlapping with sad_ls being copied. However, a Co-variable granularity incremental checkpoint stores only (the bytestring of) sad_ls.
(2) Incremental Checkout: The analyst undoes the mapping in Cell 4 due to poor results. Returning to Cell 3’s state by (completely) loading a memory snapshot is slow as it requires reloading the corpus. However, by identifying that the states of Cells 3 and 4 differ only by sad_ls, we can only load the value of sad_ls from Cell 3 (red) to replace value of sad_ls from Cell 4 (blue) to perform incremental checkout without touching the rest of the session state.
@@ 3.1. Kishu Components
Kishu (see LABEL:fig:system_overview) interacts with notebook sessions through inserting non-intrusive hooks. These hooks allow Kishu to transparently (1) monitor the kernel namespace to track session state evolution, (2) write data in the session state to storage for incremental checkpointing, and (3) alter the session state when checkout is requested.
@@
The State Loader restores to a session state upon requested checkout. It first identifies the difference between the current session state (i.e., existing items in the namespace) and the target state according to the Checkpoint Graph, then loads only the necessary data from the Checkpoint Graph for replacing the Co-variables that need updating (§ 5.2).
@@ 3.2. Kishu Workflow
This section describes Kishu’s operations during a notebook workflow. Users will interact with Kishu using the in-Jupyter Command Palette (LABEL:fig:system_overview), such as to attach it to a started notebook session; [^3] Kishu’s interface is similar to that of ElasticNotebook; a depiction can be found in its demo paper (Li et al., 2024). However, the core functionality and techniques used are different. Kishu then monitors the namespace for state deltas to incrementally checkpoint after each cell execution (if automatic checkpointing is enabled), and perform checkout to a previous state when requested.
@@
After each cell execution, the Delta Detector uses the Patched Namespace to identify updated Co-variables and stores them in a new incremental checkpoint/node with a unique checkpoint_id on the Checkpoint Graph. The user may view the stored graph, checkpoints, and their IDs with log.
Incrementally Restoring a State
Kishu will restore a previous session state with checkout checkpoint_id. The State Restorer identifies differing Co-variables (i.e., to restore) between the current and target states according to the Checkpoint Graph, then accordingly loads only the necessary data for restoration. If necessary, the Data Restorer reconstructs data that is missing (i.e., failed to serialize) or failed to load via fallback recomputation.
@@ 4.1. Co-variables
In this section, we introduce the Co-variable—Kishu’s granularity for efficient incremental checkpointing and checkout.
Preliminary: Variables, Objects and Reachability
In Python and the Jupyter Notebook ecosystem, variables and objects are 2 distinct concepts: A variable is a named entity from which various objects are reachable—for an example list ls=[1,2,3], the list name (ls) is a variable and each list element (1, 2, 3) is an object. We define reachability reference-wise, i.e., object y is reachable from variable x if y can be accessed from x through a chain of references. Some common reachability patterns include subscripting (e.g., y = x[0]), class member (e.g., y = x.attr), and attribution (e.g., y = x.__dict__). Given our distinction between variables and objects and reachability definition, we now define Co-variables as follows:
@@
For example, the Co-variable {ser,obj} in LABEL:fig:checkpoint_variable_update is modified node-wise with an in-place update “ser.replace’’ (bottom), and edge-wise with a re-assignment “obj.foo = ser[2](bottom-right). Co-variables can also be created and deleted via split and merge (right): the Co-variable {obj,ser} is deleted via a split as obj and ser longer share references, and the Co-variable {obj, st} is created through a merge. We collectively refer to Co-variable modifications, creations, and deletions as updates; the set of Co-variables updated by a cell execution is the execution’s state delta.
@@ 4.2. Accurate State Delta Detection
Kishu uses VarGraphs to detect Co-variable membership and updates. The VarGraph is a graph structure constructed from each variable in the namespace that captures its reachable objects. An example is shown in LABEL:fig:checkpoint_id_graph: each node in a variable’s VarGraph corresponds to a reachable object, containing the (1) object type, (2) memory address, and one of (3) pointers to other reachable objects (i.e., children) for non-primitive objects, or (4) value for primitive objects. For example, the node for the list reachable from ser contains 3 child pointers to the 3 nodes for strings ’a’, ’b’, and ’c’, and the node for string ’b’ holds its value ’b’.  [^4] The VarGraph is inspired by ElasticNotebook’s ID graph (Li et al., 2023a) which captures reachable objects’ memory addresses; VarGraphs uniquely contain datatypes and primitive values for additional robustness (e.g., detecting a different primitive in the same address).
@@
As Kishu constructs VarGraphs following object rechability, it detects Co-variable updates with no false negatives (verified empirically in § 7.2.1). However, Kishu’s update detection is conservative: there may be false positives if objects are dynamically generated (e.g., datatype objects) with a different memory address each time during VarGraph construction/object traversal, or cannot be traversed into (i.e., lacking referencing instructions, e.g., generators (Foundation, 2023b), which Kishu assumes to be updated on access).
@@ 4.3. Efficient State Delta Detection
Co-variable accesses indicates possible updates (e.g., via a member function call ser.replace). Kishu patches the accessor, setter, and deletion methods of the global namespace (LABEL:fig:checkpoint_detect_fast) to capture variable (hence Co-variable) accesses, which helps identify the possibly updated Co-variables: if the members of a Co-variable $(x_{1},...,x_{i})$ overlaps with the cell execution’s accessed variables, then it may have been updated (e.g., {ser,obj} in LABEL:fig:checkpoint_detect_fast): Kishu will verify the update by (1) re-generating VarGraphs for its member variables, (2) comparing the VarGraphs with those before the cell execution to identify modifications, and (3) intersecting the VarGraphs amongst variables of accessed Co-variables to identify merges and splits. Otherwise, the Co-variable surely wasn’t updated and Kishu skips its check for this cell execution (e.g., {df} in LABEL:fig:checkpoint_detect_fast, greyed out).
@@
Suppose not, i.e., Co-variable $\mathcal{X}$ has empty intersection with the accessed variables and was updated. Then, the Co-variable must have updated through another variable $y$ that was not part of the Co-variable $x_{1},...,x_{i}$ before the start of the cell execution. Due to Co-variables’ self-containment (§ 4.1), the user cannot possibly access objects reachable from $x_{1},...,x_{i}$ via $y$ during the cell execution without creating a reference by using one of $x_{1},...,x_{i}$ first (e.g., $y.foo$ = $x\_i$), but doing so violates our assumption. ∎
@@
As Kishu patches the global namespace of the notebook session, it is impossible for the user to use variables from within the notebook (i.e., to modify objects) undetected. Hence, Kishu will not misidentify Co-variables possibly updated via references.
The user may still use non-referencing methods to update data such as with direct C-pointer-based modifications, but these cases are rare in notebooks (found in 0 out of 60 surveyed notebooks (Li et al., 2023a)). Some libraries, such as NumPy (Team, 2024a), do perform memory-based updates (e.g., via slicing). However, the objects are supported by Kishu as their memory-based updates are still invoked via referencing (e.g., arr[0,1] += 1, empirically verified in § 7.2.1).

### v4

@@ Abstract.
PVLDB Reference Format:
Kishu: Time-Traveling for Computational Notebooks. PVLDB, 18(4): 970 - 985, 2024.
doi:10.14778/3717755.3717759 This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 4 ISSN 2150-8097.
doi:10.14778/3717755.3717759
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at %leave␣empty␣if␣no␣availability␣url␣should␣be␣sethttps://github.com/illinoisdata/kishu-vldb.
@@ 2.2. Characteristics of Notebook Workloads
Notebook workflows often follow a sequential [^2] Existing notebook systems, e.g., Jupyter, do not support concurrent cell executions. loop of writing and running cell code, then observing output (Rule et al., 2018). The notebook kernel can be inactive mid-loop—$\sim$10 seconds of think time (Eichmann et al., 2020) when users decide what cells to write/run next. Notebook systems use think time for computations (e.g., pre-loading data (Xin et al., 2021)). Kishu can also leverage think time for checkpointing.
@@
Python functions in notebooks can have complex access patterns (e.g., non-parameter variable accesses) that make them difficult and/or costly to analyze. Some works use rule-based approaches to hard-code effects of common functions (e.g., print not modifying the namespace) (Macke et al., 2020). however, data scientists also often import custom functions into notebooks (Rule et al., 2018); hence, an efficient approach that handles arbitrary functions is desirable.
@@
For fast iteration, users often run incremental cells each consisting of few lines of code and accessing few variables (Kery et al., 2018): we observe this in on of our test notebooks, Sklearn (LABEL:fig:background_access_objs), where 40/44 cells access ¡10% of state data. For these workloads, Kishu’s incremental checkpointing can offer larger benefits.
@@
Notebook providers often limit session memory consumption (Google, 2024; Kaggle, 2024). Hence, while creating new data, users also modify/delete existing data to conserve memory (e.g., after they have been used to generate relevant figures): in Sklearn (LABEL:fig:background_append_update_objs), we observe a 45:55 split between created and modified data. Kishu must efficiently undo these modifications during checkout.
@@ 2.3. Enabling Time Traveling
Tools such as CRIU (CRIU, 2023) snapshot notebook process memory to save session state data. (Incremental Checkpointing) Subsequent snapshots can be made incrementally w.r.t. prior snapshots storing only dirty memory pages. However, memory-page granularity is too coarse for notebook workloads as Python data structures (e.g., lists) can be fragmented, on which operations (e.g., in-place list sorting) can cause multiple dirty pages and high checkpoint costs (§ 7.3). (Complete Checkout) Memory snapshots must be entirely loaded for state restoration and require killing the existing notebook process (to avoid PID conflicts) before restoration, which is not seamless and incurs high data loading costs (§ 7.5). OS-level snapshotting is also limited to single processes, hence fail on notebooks utilizing multiple/remote processors (§ 7.3).
Application-level Time-Traveling (Ours)
We use (1) application-level information to detect state deltas at a finer granularity versus memory snapshots and (2) existing data in the kernel for incremental checkout to address the drawbacks of OS-level snapshotting: (Incremental and Generalized Checkpointing) We track state deltas at a more logical Co-variable granularity (described in § 2.4) by tracking in-notebook references for incremental checkpointing. For generalizability, we use objects’ reductions as storage instructions to checkpoint multiprocessing and off-CPU workloads (§ 6.1). (Incremental Checkout) As we can directly access the notebook kernel at the application level, we can incrementally checkout by computing the difference between the current and target states (e.g., via versioning (Bernstein and Goodman, 1983)) and updating only differing kernel data (§ 5.2).
@@ 2.4. Tracking State Delta for Time Traveling
Some existing notebook systems (Macke et al., 2020; Koop and Patel, 2017; Li et al., 2023a) track state deltas via provenance-based code analysis (e.g., via ASTs (Foundation, 2023a)) at variable-level granularity. Pure static analysis requires conservativeness on identifying changed data w.r.t. control flows (e.g., if(x<1)) and external function calls (§ 2.2), causing false positives (e.g., assuming an untaken branch as taken) and large deltas; hence, these systems augment static analysis with varying levels of live instrumentation at cell runtime (e.g., resolving x’s value when evaluating if(x<1)(Macke et al., 2020), which can result in high overhead (e.g., repeated resolutions in loops, § 7.6).
@@
To avoid potential inefficiencies of provenance tracking’s runtime resolutions, we propose performing live object comparison (i.e., comparing data pre/post-execution) only between cell executions to track per-cell updates. Our intuition is that while doing so at variable-level granularity (like existing provenance trackers) by comparing all state objects can be costly (§ 7.6), it is also unnecessary as storing/loading individual variables (e.g., with variable-level KV-stores (shove, 2024; 0xnurl, 2024)) for checkpoint/checkout risks breaking shared references (Li et al., 2023a). We hence track updates at a coarser Co-variable level—connected components of objects (w.r.t. pointer references): we can reason from access patterns which Co-variables were possibly/surely not updated by each cell to limit object comparisons, and correctly store/load Co-variables as if they are independent data tables. LABEL:fig:background_variable_blob depicts our idea: {ser,obj} is a Co-variable (red) as objects reachable from ser and obj overlap (&ser[1]=&obj.foo). {df} is another Co-variable (blue), and users cannot reach objects under df from objects under ser/obj via references. Notably, Co-variables are the minimum granularity for saving/loading state data without risking breaking shared references. § 4 formally describes Co-variables and how we correctly and efficiently capture state delta at this granularity.
@@
A data analyst performs text mining by loading the corpus (Cell 1), defining category lists (Cell 2), and sorting texts by sentiment into the lists (Cell 3). They checkpoint the state after each cell execution. Incremental Checkpointing: The analyst tests a mapping to clean the contained text in the lists on sad_ls (Cell 4, blue). Due to its interleaved construction (with other lists), sad_ls is fragmented; incrementally checkpointing at memory page granularity for Cell 4 (w.r.t. Cell 3) copies all pages overlapping with sad_ls. However, a Co-variable granularity incremental checkpoint stores only (the bytestring of) sad_ls. Incremental Checkout: The analyst undoes Cell 4’s mapping due to poor results. Returning to Cell 3’s state by (completely) loading a memory snapshot is slow as it also reloads the corpus. However, noting that Cell 3’s and 4’s states differ only by sad_ls, we can only load Cell 3’s sad_ls (red) to replace Cell 4’s sad_ls (blue) to incrementally checkout without touching the rest of the state.
@@ 3.1. Kishu Components
Kishu (LABEL:fig:system_overview) interacts with notebook sessions via non-intrusive hooks, which allow Kishu to transparently (1) monitor the namespace to track session state evolution, (2) write state data to storage for checkpointing, and (3) alter the state on requested checkouts.
@@
The State Loader restores to a session state upon requested checkout. It first identifies the difference between the current (i.e., existing items in the namespace) and target states via the Checkpoint Graph, then loads necessary data from the Checkpoint Graph to replace Co-variables that need updating (§ 5.2).
@@ 3.2. Kishu Workflow
This section covers Kishu’s operations during a notebook workflow. Users interact with Kishu with the in-Jupyter Command Palette (LABEL:fig:system_overview), such as to attach it to a new notebook session; [^3] A depiction of Kishu’s interface can be found in our prior work’s demo paper (Li et al., 2024). Kishu then monitors the namespace for state deltas, checkpoint after each cell execution, and perform on-demand checkout to previous states.
@@
After each cell execution, the Delta Detector uses the Patched Namespace to identify updated Co-variables and stores them in a new incremental checkpoint/node with a unique ckpt_id on the Checkpoint Graph. The user may view the stored graph, checkpoints, and their IDs with log.
Incremental Checkout
Kishu will restore a previous session state with checkout ckpt_id. The State Restorer identifies differing Co-variables between the current and target states according to the Checkpoint Graph, then accordingly loads only the necessary data for restoration. If necessary, the Data Restorer reconstructs data that is missing or failed to load via fallback recomputation.
@@ 4.1. Co-variables
This section introduces Kishu’s definition of the Co-variable.
Preliminary
In Python and the Jupyter Notebook ecosystem, variables and objects are 2 distinct concepts: A variable is a named entity from which various objects are reachable—for an example list ls=[1,2,3], the list name (ls) is a variable and each list element (1, 2, 3) is an object. We define reachability reference-wise, i.e., object y is reachable from variable x if y can be accessed from x through a chain of references. Some common reachability patterns include subscripting (e.g., y=x[0]) and class member (e.g., y=x.attr). Given this distinction between variables and objects and reachability definition, we now define Co-variables as follows:
@@
For example, the Co-variable {ser,obj} in LABEL:fig:checkpoint_variable_update is modified node-wise with an in-place update “ser.replace’’ (bottom), and edge-wise with a re-assignment “obj.foo=ser[2](bottom-right). Co-variables can be created and deleted via split and merge (right): {obj,ser} is deleted via a split as obj and ser longer share references, and {obj,st} is created via a merge. We collectively refer to Co-variable modifications, creations, and deletions as updates; the Co-variables updated by a cell execution form its state delta.
@@ 4.2. Accurate State Delta Detection
Kishu detects Co-variable membership and updates with VarGraphs—a graph structure constructed on each variable that captures its reachable objects in the namespace. LABEL:fig:checkpoint_id_graph shows an example: each node in a variable’s VarGraph corresponds to a reachable object, containing the (1) type, (2) memory address, and one of (3) pointers to other reachable objects (i.e., children) for non-primitives, or (4) value for primitives. For example, the node for list contains 3 child pointers to the 3 nodes for strings ’a’, ’b’, and ’c’, and the node for string ’b’ holds its value ’b’.  [^4] The VarGraph is inspired by ElasticNotebook’s ID graph (Li et al., 2023a) which captures reachable objects’ memory addresses; VarGraphs uniquely contain datatypes and primitive values for additional robustness (e.g., detecting a different primitive in the same address).
@@
As Kishu constructs VarGraphs following object rechability, it detects Co-variable updates with no false negatives (empirically verified in § 7.2.1). However, Kishu’s update detection is conservative: there may be false positives if objects are dynamically generated (e.g., datatype objects) with a different memory address each time during VarGraph construction/object traversal, or cannot be traversed into (i.e., lacking referencing instructions, e.g., generators (Foundation, 2023b), which Kishu assumes to be updated on access).
@@ 4.3. Efficient State Delta Detection
Co-variable accesses indicates possible updates (e.g., via a member function call ser.replace). Kishu patches the accessor, setter, and deletion methods of the global namespace (LABEL:fig:checkpoint_detect_fast) to capture variable (hence Co-variable) accesses, which helps identify the possibly updated Co-variables: if the members of a Co-variable $\mathcal{X}$ overlaps with the cell execution’s accessed variables, then it may have been updated (e.g., {ser,obj} in LABEL:fig:checkpoint_detect_fast): Kishu will verify the update by (1) re-generating VarGraphs for its member variables, (2) comparing the VarGraphs with those before the cell execution to identify modifications, and (3) intersecting the VarGraphs amongst variables of accessed Co-variables to identify merges and splits. Otherwise, the Co-variable surely wasn’t updated and Kishu skips its check for this cell execution (e.g., {df} in LABEL:fig:checkpoint_detect_fast, greyed out).
@@
Suppose not, i.e., Co-variable $\mathcal{X}$ does not intersect the accessed variables and was updated. Then, $\mathcal{X}$ must have been updated through via variable $y$ that was not part of $\mathcal{X}$ before the start of the cell execution. Due to Co-variables’ self-containment (§ 4.1), objects reachable from $\mathcal{X}$ cannot possibly be accesed via $y$ during the cell execution without creating a reference by using one of $x_{1},...,x_{i}$ first (e.g., $y.foo$ = $x\_i$), but doing so violates our assumption. ∎
@@
As Kishu patches the notebook session’s global namespace, it is impossible for users to use variables from within the notebook (e.g., to modify objects) undetected. Hence, Kishu will not misidentify Co-variables possibly updated via references. Users may still use non-referencing methods to update data such as C-pointer-based modifications, but these cases are rare in notebooks (found in 0/60 surveyed notebooks (Li et al., 2023a)) Some libraries do perform memory-based updates (e.g. NumPy’s slicing (Team, 2024a)). However, the objects are supported by Kishu as these updates are still invoked via referencing (e.g., arr[0,1] += 1, empirically verified in § 7.2.1).